# Multi-Group Proportional Representation

### v1

@@ 6 Concluding Remarks and Limitations
Our work has limitations that offer opportunities for future research. First, while MPR is designed to handle a large number of groups, the computational complexity of the proposed cutting-plane algorithm in the case of a general function class $\mathcal{C}$ can be a concern for high-dimensional feature spaces. Developing more efficient algorithms or approximation techniques could help address this issue. Second, while we focused on retrieval tasks, adapting MPR to other domains, such as ranking, recommendation systems, and image generation, could broaden its impact and applicability. Future research could also investigate the long-term societal impact of enforcing MPR in real-world retrieval systems to understand its potential benefits and risks. Finally, MPR is inherently limited by the curated dataset. If this dataset is biased against a given group, this bias will be propagated by MAPR. The creation and selection of curated datasets must carefully account for stakeholder interest and target representation statistics.
@@ 7 Acknowledgements
Supplementary material to the paper Multi-Group Proportional Representation
In this supplement to the paper, we present the proofs in Appendix A. In particular, we prove Proposition 1, Proposition 2, and Proposition 4. In Appendix B, we discuss the MPR-aware retrieval algorithm, its convergence aspects, and the MPR-aware retrieval for linear functions and functions belonging to an RKHS. In Appendix C, we describe the algorithms, provide further retrieval statistics, and a brief discussion about proportional Representation in political sciences. In particular, after describing the algorithms CLIP-clip Wang et al. [18], CLIP-debias Berg et al. [20], Post-Hoc Bias Mitigation ([29] and PATHS (Srinivasan et al. [27]), we include additional plots for three datasets for decision tree models in Fig. 10 for multilayer perceptron in Fig. 11. In Appendix E, we provide experiment Setting and details. Finally, in Appendix D, discuss additional experimental results and compare MAPR  against $k$-NN, MMR, ClipClip, PBM, and DebaisCLIP. In particular, we include full tables of intersectional retrieval performance on UTKFace.
Appendix A Proofs from Section 3
@@
We start by proving Proposition 1 from Section 3 which bounds the error in MPR estimates when $Q$ is approximated via a curated dataset $\mathcal{D}_{C}$. We do so by using tools from standard Rademacher complexity results in statistical learning theory, e.g., [70]. See also [71, Chapter 4]. We restate the theorems here to facilitate readability. We start with a lemma about the difference of suprema.
@@
$\sup_{c\in\mathcal{C}}\mathbb{E}_{P}[c(X)]-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)]$
$=\sup_{c\in\mathcal{C}}\left\{\mathbb{E}_{P}[c(X)]-\mathbb{E}_{P^{\prime}}[c(X)]+\mathbb{E}_{P^{\prime}}[c(X)]\right\}-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)]$
$\leq\sup_{c\in\mathcal{C}}\left\{\mathbb{E}_{P}[c(X)]-\mathbb{E}_{P^{\prime}}[c(X)]\right\}+\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)]-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)]$
@@
Proposition 5 (Generalization Gap of MPR).
Let $\{\boldsymbol{x}^{r}_{i}\}_{i=1}^{k}$ be retrieved samples $\mathcal{R}$ and $\{\boldsymbol{x}^{c}_{i}\}_{i=1}^{m}$ be i.i.d. samples from a curated distribution $Q$ used to generate a curation dataset $\mathcal{D}_{C}$. If $\mathcal{C}=\{c:\mathcal{X}\rightarrow\{-1,1\}\}$ and ${\mathcal{R}}_{m}(\mathcal{C})$ is the Rademacher complexity of $\mathcal{C}$ then, for all $\delta>0$
$\left|\mathrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\mathrm{MPR}(\mathcal{C},\mathcal{R},Q)\right|\leq{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{8m}}$ (12)
with probability at least $1-\delta$.
@@
$=\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|$ (14)
$\leq\sup_{c\in\mathcal{C}}\left\{\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|\right\}$ (15)
$\leq\sup_{c\in\mathcal{C}}\left|\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|\right|$ (16)
$\leq\sup_{c\in\mathcal{C}}\left|\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})-\mathbb{E}_{Q}[c(X)]\right|$ (17)
@@
$\Pr\left[\left|\textrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|\right|\geq 2{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{2m}}\right]$ (19)
$\leq\Pr\left[\sup_{c\in\mathcal{C}}\left|\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})-\mathbb{E}_{Q}[c(X)]\right|\geq 2{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{2m}}\right]$ (20)
@@
Proposition 6 (Query Budget Guarantee).
Denote queries to retrieve samples by $\mathcal{Q}=\{q_{1},...,q_{M}\}$ where $M\in\mathbb{N}$ is the number of queries. Let $\mathsf{VC}(\mathcal{C})$ denote the VC-dimension of the class $\mathcal{C}$ with image in $\{-1,1\}$. Then, for $\epsilon>0$ if the curation dataset contains $m\in\mathbb{N}$ samples from the curated data distribution $Q$ such that
$m\geq\frac{128\texttt{VC}(\mathcal{C})}{\epsilon^{2}}+\frac{2\log\left(\frac{2M}{\delta}\right)}{\epsilon^{2}}$ (22)
then we have with probability at least $1-\delta$ that
$\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|\leq\textrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})+\epsilon$ (23)
@@
$\Pr\left[\sup_{q\in\mathcal{Q}}\left|\mathrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\mathrm{MPR}(\mathcal{C},\mathcal{R},Q)\right|\geq{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta^{*}}\right)}{8m}}\right]$
$\leq\sum_{q\in\mathcal{Q}}\Pr\left[\left|\mathrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\mathrm{MPR}(\mathcal{C},\mathcal{R},Q)\right|\geq{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta^{*}}\right)}{8m}}\right]$
@@
$\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i}(q))-\mathbb{E}_{Q}[c(X)]\right|$ $\leq\textrm{MPR}(\mathcal{C},\mathcal{R}(q),\mathcal{D}_{C})+{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2M}{\delta}\right)}{8m}}$ (26)
@@
$m\geq\frac{32\texttt{VC}(\mathcal{C})}{\epsilon^{2}}+\frac{\log\left(\frac{2M}{\delta}\right)}{2\epsilon^{2}}$ (28)
@@
$\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i}(q))-\mathbb{E}_{Q}[c(X)]\right|\leq\textrm{MPR}(\mathcal{C},\mathcal{R}(q),\mathcal{D}_{C})+\epsilon$ (29)
@@
Proposition 7.
Let $\mathcal{C}^{\prime}\triangleq\{c\in\mathcal{C}\mid\sum_{i=1}^{m+n}c(\boldsymbol{x}_{i})^{2}=\frac{mk}{m+k}\}$ where $\mathcal{C}$ is closed under scalar multiplication. Let $c(\mathbf{X})=[c(\mathbf{x_{1}}),...,c(\mathbf{x_{n+m}})]$. Then
$0\leq\mathrm{MPR}(\mathcal{C}^{\prime},\mathcal{R},\mathcal{D}_{C})\leq 1,$ (30)
and for any
let $\hat{c}:\mathcal{X}\rightarrow\mathbb{R}$ be defined as $\hat{c}(\mathbf{x})=\frac{\sqrt{C}}{\|c^{*}(\mathbf{X})\|_{2}}c^{*}(\mathbf{x})$. Then, we have that
@@
We begin proving the first statement. As a shorthand, let $c(\mathbf{X})=[c(\mathbf{x_{1}}),...,c(\mathbf{x_{n+m}})]$. The Cauchy-Schwarz inequality gives,
@@
assuming $\|c(\mathbf{X})\|_{2}<\infty$. Plugging this back into Equation (36), the MSE problem becomes
$\operatorname*{arg\,inf}_{c\in\mathcal{C}}\sum_{i=1}^{n+m}(\frac{\sum_{i=1}^{n+m}c(\mathbf{x}_{i})\widetilde{a}_{i}}{\|c(\mathbf{X})\|_{2}^{2}}c(\boldsymbol{x}_{i})-\widetilde{a}_{i})^{2}$ (38)
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}\sum_{i=1}^{n+m}\left(\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{4}}c(\boldsymbol{x}_{i})^{2}-2\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle}{\|c(\mathbf{X})\|_{2}^{2}}c(\boldsymbol{x}_{i})\widetilde{a}_{i}+\widetilde{a}_{i}^{2}\right)$ (39)
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}\left(\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{4}}\|c(\mathbf{X})\|_{2}^{2}-2\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle}{\|c(\mathbf{X})\|_{2}^{2}}\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle+\sum_{i=1}^{n+m}\widetilde{a}_{i}^{2}\right)$ (40)
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}-\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{2}}+\sum_{i=1}^{n+m}\widetilde{a}_{i}^{2}$ (41)
@@
$\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{2}}=\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{|\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle|^{2}}{\|c(\mathbf{X})\|_{2}^{2}}$ (42)
@@
$c^{*}\in\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{|\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle|^{2}}{\|c(\mathbf{X})\|_{2}^{2}},$ (44)
@@
Now, we establish Proposition 4, restated below for convenience. It says that MPR enjoys a closed-form expression whenever the class $\mathcal{C}$ is given by linear functions. This proposition allows us to simplify MAPR  since, in this case, the optimization problem can be cast as a quadratic program.
Proposition 8.
where $\mathbf{X}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\intercal}$ is the SVD of $\mathbf{X}$ and $\mathbf{U}_{l}\in\mathbb{R}^{(m+n)\times l}$ are the left singular vectors in $\mathbf{U}$ corresponding to the top-$l$ largest singular values.
The closed-form expression for MPR in (9) allows MPR-constrained retrieval to be approximated by a quadratic program, whose proof can be found in Appendix B. Moreover, the above proposition can be directly adapted to linear functions defined only for embeddings or a subset of features of retrieved items.
@@
subject to $\sum_{i=1}^{m+n}[w^{\intercal}x]^{2}=\frac{mk}{m+k},$ (49)
@@
subject to $\|\mathbf{X}\textbf{w}\|_{2}^{2}=\frac{mk}{m+k},$ (51)
@@
subject to $\|\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}\|_{2}^{2}=\frac{mk}{m+k}.$ (53)
Now, by a change of variables $\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}=\widetilde{\textbf{w}}$ and $\mathbf{U}^{\intercal}\mathbf{\widetilde{a}}=\mathbf{z}$
@@
subject to $\|\widetilde{\textbf{w}}\|_{2}^{2}=\frac{mk}{m+k},$ (55)
$\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}=\widetilde{\textbf{w}}.$ (56)
@@
$\frac{\partial L}{\partial\widetilde{\textbf{w}}}=\mathbf{z}-\lambda\widetilde{\textbf{w}}+\boldsymbol{\mu}=0,$ $\ \frac{\partial L}{\partial\textbf{w}}=-\mathbf{V}\mathbf{\Sigma}\boldsymbol{\mu}=0,$ (58)
$\frac{\partial L}{\partial\lambda}=\frac{1}{2}\left(\|\widetilde{\textbf{w}}\|_{2}^{2}-\frac{mk}{m+k}\right)=0,$ $\ \frac{\partial L}{\partial\boldsymbol{\mu}}=\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}-\widetilde{\textbf{w}}=0.$ (59)
From the second equation, we have $\mathbf{V}\mathbf{\Sigma}\boldsymbol{\mu}=0$. Assuming $\mathbf{V}$ and $\mathbf{\Sigma}$ are full rank, this implies $\boldsymbol{\mu}=0$. Substituting this into the first equation, we get:
@@
Using the third equation and the constraint $\|\widetilde{\textbf{w}}\|_{2}^{2}=\frac{mk}{m+k}$, we can solve for $\lambda$:

### v2

@@ 7 Acknowledgements
Supplementary material to the paper Multi-Group Proportional Representation in Retrieval
In this supplement to the paper, we begin in Section A and Section B by noting some ethical considerations and limitations for our work, respectively. Then, we present our proofs of Propositions 1, 3, 2, and 4 in Appendix C. In Appendix D, we discuss the MOPR algorithm, its convergence aspects, and MOPR for linear functions and functions belonging to an RKHS. We also compare the linear relaxation to the integer program optimization problem in Equation 9 empirically and similarly compare the result of MOPR to directly solving the closed form linear program in the case of linear models. In Appendix E, we describe the baselines reported in this paper and provide further retrieval statistics. In particular, we describe the algorithms CLIP-clip Wang et al. [18], CLIP-debias Berg et al. [20], Post-Hoc Bias Mitigation ([29] and PATHS [27]. In F, we provide a brief discussion about proportional representation in political sciences, give a few historical comments about social choice theory, and discuss the contribution given by [37]. Then, in Appendix G, we include an additional table of results for equal representation with the other baseline methods (Table 5). We also include plots for additional queries for the linear regression model (Figs. 8, 9, and 10) for as well as two additional models, 1) decision tree models in Figs. 11, 12, and 13 and for 2) multilayer perceptrons in Figs. 14, 15, and 16. We provide additional details on the experiment setting in Appendix H. Finally, we include some example sets of retrieved images for qualitative comparison.
Appendix A Ethics considerations
It is important to consider the long-term societal impact of enforcing MPR in real-world retrieval systems to understand its potential benefits and risks. First, applying MPR, such as with MOPR in deployment settings, could result in “ethics-washing"–where a company claims their system is fair because they measured MPR when, in fact, there are still representational harms–or could provide users and developers with a false sense of fairness. Second, MPR is just one of many ways of analyzing representation, and as a statistical metric, it fails to consider aspects such as human perceptions of representation. There are legal and regulatory risks with overreliance on a single metric for fairness, especially if this metric is used to inform policy and decision-making. As mentioned previously, any biases present in the curation dataset will propagate to the results and downstream uses of MPR; furthermore, the choice of what dataset and what attributes to consider is an inherently social and political choice made by humans, subject to its own limitations and biases. Finally, the introduction of MPR in real-world retrieval systems could introduce new stereotypes or erase identities if not conducted properly.
Appendix B Limitations
Our aggregate results (e.g., Table 1) should be approached with caution, as each MPR-similarity curve is sampled from a different distribution of data conditioned on a query, limiting our ability to draw statistical conclusions. This does not preclude the existence of a query where a competing method may achieve an MPR-similarity trade-off point not achieved by MOPR. However, we note that MOPR is specifically optimized to promote MPR while preserving similarity, explaining its favorable performance. In all instances observed by the authors, MOPR Pareto dominates competing benchmarks and was the most successful method in achieving low values of MPR while maintaining high levels of similarity with a given query. Second, while MPR is designed to handle a large number of groups, the computational complexity of the proposed cutting-plane algorithm in the case of a general function class $\mathcal{C}$ can be a concern for high-dimensional feature spaces. Developing more efficient algorithms or approximation techniques could help address this issue. Finally, MPR is inherently limited by the curated dataset. If this dataset is biased against a given group, this bias will be propagated by MOPR. The creation and selection of curated datasets must carefully account for stakeholder interest and target representation statistics. Additionally, we note that not every combination of intersectional identities (such as “white male" as opposed to “black female") may be considered intersectional from a sociological perspective, and furthermore, these datasets do not contain multiracial labels. We do not presume to address representational issues in ML datasets, and instead, our experiments aim to demonstrate the flexibility and scalability of the MPR metric in handling complex intersectional groups. We hope that MPR will only prove more useful as datasets become more representative.
Appendix C Proofs from Section 3
@@
We start by proving Proposition 1 from Section 3 which bounds the error in MPR estimates when $Q$ is approximated via a curated dataset $\mathcal{D}_{C}$. We do so by using tools from standard Rademacher complexity results in statistical learning theory, e.g., [72]. See also [73, Chapter 4]. We restate the theorems here to facilitate readability. We start with a lemma about the difference of suprema.
@@
$\sup_{c\in\mathcal{C}}\mathbb{E}_{P}[c(X)]-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)],$
$=\sup_{c\in\mathcal{C}}\left\{\mathbb{E}_{P}[c(X)]-\mathbb{E}_{P^{\prime}}[c(X)]+\mathbb{E}_{P^{\prime}}[c(X)]\right\}-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)],$
$\leq\sup_{c\in\mathcal{C}}\left\{\mathbb{E}_{P}[c(X)]-\mathbb{E}_{P^{\prime}}[c(X)]\right\}+\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)]-\sup_{c\in\mathcal{C}}\mathbb{E}_{P^{\prime}}[c(X)],$
@@
See 1
@@
$=\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\mathbb{E}_{Q}[c(X)]\right|,$ (12)
$\leq\sup_{c\in\mathcal{C}}\left\{\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\mathbb{E}_{Q}[c(X)]\right|\right\},$ (13)
$\leq\sup_{c\in\mathcal{C}}\left|\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|-\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i})-\mathbb{E}_{Q}[c(X)]\right|\right|,$ (14)
$\leq\sup_{c\in\mathcal{C}}\left|\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})-\mathbb{E}_{Q}[c(X)]\right|.$ (15)
@@
$\Pr\left[\left|\textrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(x_{i})-\mathbb{E}_{Q}[c(X)]\right|\right|\geq 2{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{2m}}\right],$ (17)
$\leq\Pr\left[\sup_{c\in\mathcal{C}}\left|\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})-\mathbb{E}_{Q}[c(X)]\right|\geq 2{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta}\right)}{2m}}\right],$ (18)
@@
See 2
@@
$\Pr\left[\sup_{q\in\mathcal{Q}}\left|\mathrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\mathrm{MPR}(\mathcal{C},\mathcal{R},Q)\right|\geq{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta^{*}}\right)}{8m}}\right],$
$\leq\sum_{q\in\mathcal{Q}}\Pr\left[\left|\mathrm{MPR}(\mathcal{C},\mathcal{R},\mathcal{D}_{C})-\mathrm{MPR}(\mathcal{C},\mathcal{R},Q)\right|\geq{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2}{\delta^{*}}\right)}{8m}}\right],$
@@
$\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i}(q))-\mathbb{E}_{Q}[c(X)]\right|$ $\leq\textrm{MPR}(\mathcal{C},\mathcal{R}(q),\mathcal{D}_{C})+{\mathcal{R}}_{m}(\mathcal{C})+\sqrt{\frac{\log\left(\frac{2M}{\delta}\right)}{8m}},$ (22)
@@
$m\geq\frac{32\texttt{VC}(\mathcal{C})}{\epsilon^{2}}+\frac{\log\left(\frac{2M}{\delta}\right)}{2\epsilon^{2}},$ (24)
@@
$\sup_{c\in\mathcal{C}}\left|\frac{1}{k}\sum_{i=1}^{k}c(\boldsymbol{x}^{r}_{i}(q))-\mathbb{E}_{Q}[c(X)]\right|\leq\textrm{MPR}(\mathcal{C},\mathcal{R}(q),\mathcal{D}_{C})+\epsilon.$ (25)
@@
See 3
@@
We begin proving $0\leq\mathrm{MPR}(\mathcal{C}^{\prime},\mathcal{R},\mathcal{D}_{C})\leq 1,$. As a shorthand, let $c(\mathbf{X})=[c(\mathbf{x_{1}}),...,c(\mathbf{x_{n+m}})]$. The Cauchy-Schwarz inequality gives,
@@
as $\|c(\mathbf{X})\|_{2}<\infty$. Plugging this back into Equation (29), the MSE problem becomes
$\operatorname*{arg\,inf}_{c\in\mathcal{C}}\sum_{i=1}^{n+m}(\frac{\sum_{i=1}^{n+m}c(\mathbf{x}_{i})\widetilde{a}_{i}}{\|c(\mathbf{X})\|_{2}^{2}}c(\boldsymbol{x}_{i})-\widetilde{a}_{i})^{2},$
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}\sum_{i=1}^{n+m}\left(\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{4}}c(\boldsymbol{x}_{i})^{2}-2\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle}{\|c(\mathbf{X})\|_{2}^{2}}c(\boldsymbol{x}_{i})\widetilde{a}_{i}+\widetilde{a}_{i}^{2}\right),$
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}\left(\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{4}}\|c(\mathbf{X})\|_{2}^{2}-2\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle}{\|c(\mathbf{X})\|_{2}^{2}}\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle+\sum_{i=1}^{n+m}\widetilde{a}_{i}^{2}\right),$
$=\operatorname*{arg\,inf}_{c\in\mathcal{C}}-\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{2}}+\sum_{i=1}^{n+m}\widetilde{a}_{i}^{2}.$
@@
$\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle^{2}}{\|c(\mathbf{X})\|_{2}^{2}}=\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{|\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle|^{2}}{\|c(\mathbf{X})\|_{2}^{2}}.$
@@
$c^{*}\in\operatorname*{arg\,sup}_{c\in\mathcal{C}}\frac{|\langle c(\mathbf{X}),\mathbf{\widetilde{a}}\rangle|^{2}}{\|c(\mathbf{X})\|_{2}^{2}},$
@@
Now, we establish Proposition 4, restated below for convenience. It says that MPR enjoys a closed-form expression whenever the class $\mathcal{C}$ is given by linear functions. This proposition allows us to simplify MOPR  since, in this case, the optimization problem can be cast as a quadratic program.
See 4
The closed-form expression for MPR in (8) allows MPR-constrained retrieval to be approximated by a quadratic program, whose proof can be found in Appendix D. Moreover, the above proposition can be directly adapted to linear functions defined only for embeddings or a subset of features of retrieved items.
@@
subject to $\sum_{i=1}^{m+n}[w^{\intercal}x]^{2}=\frac{mk}{m+k},$
@@
subject to $\|\mathbf{X}\textbf{w}\|_{2}^{2}=\frac{mk}{m+k},$
@@
subject to $\|\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}\|_{2}^{2}=\frac{mk}{m+k}.$
Now, by a change of variables $\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}=\widetilde{\textbf{w}}$ and $\mathbf{U}_{l}^{\intercal}\mathbf{\widetilde{a}}=\mathbf{z}$
@@
subject to $\|\widetilde{\textbf{w}}\|_{2}^{2}=\frac{mk}{m+k},$
$\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}=\widetilde{\textbf{w}}.$
@@
$\frac{\partial L}{\partial\widetilde{\textbf{w}}}=\mathbf{z}-\lambda\widetilde{\textbf{w}}+\boldsymbol{\mu}=0,$ (39)
$\frac{\partial L}{\partial\textbf{w}}=-\mathbf{V}\mathbf{\Sigma}\boldsymbol{\mu}=0,$ (40)
$\frac{\partial L}{\partial\lambda}=\frac{1}{2}\left(\|\widetilde{\textbf{w}}\|_{2}^{2}-\frac{mk}{m+k}\right)=0,$ (41)
$\frac{\partial L}{\partial\boldsymbol{\mu}}=\mathbf{\Sigma}\mathbf{V}^{\intercal}\textbf{w}-\widetilde{\textbf{w}}=0.$ (42)
From (42), we have $\mathbf{V}\mathbf{\Sigma}\boldsymbol{\mu}=0$. Assuming $\mathbf{V}$ and $\mathbf{\Sigma}$ are full rank, this implies $\boldsymbol{\mu}=0$. Substituting this into (39), we get:
@@
Using (43) and the constraint $\|\widetilde{\textbf{w}}\|_{2}^{2}=\frac{mk}{m+k}$, we can solve for $\lambda$: